# Non-Linear Flow Matching for Full-Atom Peptide Design

### v1

@@ Abstract
Peptide design plays a pivotal role in therapeutic applications, yet existing AI-assisted methods often struggle to generate stable peptides with high affinity due to their inability to accurately simulate the dynamic docking process. To address this challenge, we propose NLFlow, a novel multi-manifold approach based on non-linear flow matching. Specifically, we design a polynomial-based conditional vector field to accelerate the convergence of the peptide’s position towards the target pocket, effectively capturing the temporal inconsistencies across position, rotation, torsion, and amino acid type manifolds. This enables the model to better align with the true conformational changes observed in biological docking processes. Additionally, we incorporate interaction-related information, such as polarity, to enhance the understanding of peptide-protein binding. Extensive experiments demonstrate that NLFlow outperforms existing methods in generating peptides with superior stability, affinity, and diversity, offering a fast and efficient solution for peptide design and advancing the peptide-based therapeutic development.
Machine Learning, ICML
1 Introduction
@@
Therapeutic peptides are short, single-chain proteins composed of amino acids, and their design is essential for achieving precise and stable docking with target receptor proteins (Craik et al., 2013; Fosgerau & Hoffmann, 2015). Full-atom peptide design enables the creation of such peptides, offering a robust solution for advancing therapeutic development and treating protein-related diseases (Muttenthaler et al., 2021; Petsalaki & Russell, 2008; Lee et al., 2019a). To tackle the vast and complex chemical space, recent generative models integrate protein sequence and structure data to extract rich feature representations, facilitating the exploration of sequence-structure relationships (Charoenkwan et al., 2021; Lin et al., 2023). Particularly, diffusion and flow matching models simulate peptide movement through time-dependent random pathways or vector fields (Ciemny et al., 2018; Dauparas et al., 2022; Watson et al., 2023). These models account for modalities such as displacement, rotation, and other dynamics. By imposing constraints on the peptide’s conformational degrees of freedom, they enable efficient exploration of conformational space and provide an accurate simulation of docking dynamics. Thus, they overcome the limitations of traditional experimental methods, such as high costs and long timelines (Henninot et al., 2018; Bhardwaj et al., 2016).
However, existing diffusion models neglect temporal dependencies across different modalities in the docking process, leading to generated peptides that cannot stably bind to target proteins (Bennett et al., 2023). Although these methods model multi modalities to simulate peptide movement, they treat the evolution of all these modalities as fully synchronized in time (Corso et al., 2022; Li et al., 2024; Lin et al., 2024). In reality, peptides must first move towards the target pocket before specific residue interactions can trigger subsequent conformational changes, such as backbone rotation and side-chain torsion (London et al., 2013). This temporal inconsistency during inference results in inaccurate peptide movement predictions and impairs the biologically realistic simulation of peptide dynamics.
Figure 1: Modal Temporal Sequence in Docking: Position transformations occur prior to rotational and torsional changes due to the requirement that the peptide first moves into proximity with the target pocket, allowing for residue interactions with the surrounding environment, which then drive backcone rotations and side-chain torsion adjustments.
@@
To resolve the temporal inconsistency across different manifolds, we propose NLFlow, a novel Non-Linear multi-manifold based Flow matching method. It accelerates the peptide’s coordinate convergence towards the target in a non-uniform manner, prioritizing positional alignment over rotational and torsional adjustments. This closely mirrors the actual docking process, where the peptide first moves towards the protein pocket and then fine-tunes its posture. To realize this, the core idea is to leverage polynomial interpolation to define a time-varying gradient vector, decoupling the evolution of the position manifold from the other manifolds. Specifically, the gradient magnitude of the position manifold decreases over time, accelerating the initial phase of convergence, thereby reflecting the temporal inconsistency across different manifolds. Additionally, to prevent the model from focusing solely on conformational alignment while neglecting docking interactions, interaction-related information such as polarity, charge, and hydrophilicity is incorporated, ensuring stronger binding affinity to the target protein.
@@
The NLFlow framework consists of two key components: (i) In the pre-training phase, a neural network is used to fit the gradient vector field of four manifolds: peptide position, rotation, torsion and amino acid types, learning distribution evolution of docking process; (ii) In the sampling phase, based on ordinary differential equations (ODEs) (Teschl, 2021), the gradient vector is predicted with the pre-trained model at each time step, and the peptide conformation is updated gradually over the integration of time.
NLFlow provides two key advantages: (i) Introduces a non-linear flow algorithm to address temporal inconsistency between different modalities. This leads to a more accurate simulation of peptide docking, generating more than twice as many stable peptide-protein complexes compared to the baseline; (ii) The rapid initial position convergence of the non-linear flow reduces the ineffective time spent on conformational exploration, allowing the model to gain a deeper understanding of the relationship between fine conformational changes and interaction forces, which leads to 7.13% more peptides with lower binding energies.
@@
  • We first identify the temporal inconsistencies across modalities as a critical issue in multi-modal peptide design models, which violates biological principles and hinders the generation of high-affinity peptides.
@@
  • We design a novel non-linear flow matching approach NLFlow that decouples the evolution of different manifolds, providing a more accurate simulation of dynamic conformational changes.
@@
  • Extensive experiments demonstrate NLFlow’s superior performance in both co-design and re-docking, significantly improving generated peptides’ stability and affinity, while providing an efficient method for advancing peptide-based therapeutic development.
2 Related Work
2.1 Peptide generation
@@
Peptide generation is a complex task in drug discovery, aiming to design peptides with specific biological functions, such as binding to target proteins or modulating enzymatic activity. Generative models have significantly advanced peptide design by exploring large sequence and structural spaces with greater flexibility. For example, RFDiffusion (Watson et al., 2023), designed for protein design, inspired the application of similar diffusion techniques for peptide generation. AMP-diffusion (Chen et al., 2024) utilizes the capabilities of protein large language model ESM-2 (Beal, 2015) to regenerate functional antimicrobial peptides (AMPs). MMCD (Wang et al., 2024) employs multi-modal contrastive learning in diffusion, exploiting the integration of both sequence and structural information to produce peptides with high functional relevance.
@@
Peptide design is a subtask of peptide generation, involving several key components. One important aspect is backbone design, where methods like PepFlow (Li et al., 2024) and PPFlow (Lin et al., 2024) use flow matching to simulate dynamic conformational changes and optimize peptide properties. Another aspect is side-chain packing, with methods such as RED-PPI (Luo et al., 2023) focusing on protein-protein complexes and DiffPack (Zhang et al., 2024) targeting peptide-protein interactions.
2.2 Protein–ligand docking
@@
Protein–ligand docking aims to predict the binding pose and affinity of a ligand to a target protein. Traditional docking methods often rely on rigid-body simulations and predefined scoring functions, which can struggle to handle flexible ligands and complex interactions (Kramer et al., 1999; Totrov & Abagyan, 1997). Recent years, deep learning (DL) approaches have significantly improved docking accuracy (Yang et al., 2022a). Models like AtomNet (Stafford et al., 2022) and OnionNet (Wang et al., 2021) use convolutional neural networks (CNNs) to capture complex molecular features, significantly enhancing binding affinity predictions. Further developments have incorporated graph-based models, such as GraphSite (Shi et al., 2022) and DGraphDTA (Yang et al., 2022b), which represent protein-ligand interactions as graphs to better model flexible docking. Additionally, advances in protein structure characterization, notably via AlphaFold (Evans et al., 2021), have enhanced understanding of protein flexibility, fueling the development of structure-based docking prediction methods (Wong et al., 2022; Johansson-Åkhe & Wallner, 2022). However, challenges still remain in fully capturing the dynamic and flexible nature of ligands, highlighting the need for further improvements in docking models.
3 Method
3.1 Preliminary
Conditional Flow Matching.
Conditional Flow Matching (CFM) (Lipman et al., 2022) models the transformation between a prior distribution $p=p_{0}$ and a target data distribution $q=p_{1}$ on a manifold $\mathcal{M}$ using ODEs. The transformation is described as:
$\frac{d}{dt}\varphi_{t}(x)=u_{t}(x),$ (1)
where $\varphi_{t}(x)$ represents the data distribution at time $t$, and $u_{t}(x)$ is the gradient vector field driving the flow. Since the true gradient flow $u_{t}(x)$ is unknown, it is assumed based on a hypothesis form and a known target distribution $x_{1}$. Usually a neural network is used to approximate this flow by learning a vector field $v_{t}(x)$. The loss is defined to minimize the difference between $v_{t}(x)$ and $u_{t}(x|x_{1})$:
In most current methods, the distribution $\varphi_{t}(x)$ is typically assumed to follow a simple linear interpolation between the initial distribution $x_{0}$ and target distribution $x_{1}$: $\varphi_{t}(x)=tx_{1}+(1-t)x_{0}$, where $t\in[0,1]$. The corresponding gradient vector field is $u_{t}(x)=\frac{d}{dt}\varphi_{t}(x)=x_{1}-x_{0}$. This form has been proven to correspond to the Optimal Transport (OT) solution, which provides an efficient and fast means of transforming distributions with a fixed direction and magnitude.
Problem Statement.
@@ 3.1 Preliminary
Peptide is a short sequence of amino acids, denoted by $a$, representing its type. Each amino acid is characterized by its backbone and side chain.
@@
The backbone consists of four heavy atoms $N$, $C_{\alpha}$, $C$, and $O$, with the position of $C_{\alpha}$ serving as the reference, denoted as $\mathbf{X}_{i}$ at the origin. Peptide’s orientation is modeled in two parts: the overall rigid rotation, represented by the rotation matrix $\mathbf{R}\in\operatorname{SO}(3)$, and the side-chain torsion, captures four torsion angles $\boldsymbol{\tau}=(\chi_{1},\chi_{2},\chi_{3},\chi_{4})$ of the rotatable bonds.
@@
Figure 3: Overview of the NLFlow framework
3.2 Non-Linear Flow for Position
@@
Despite being efficient and fast, the OT assumption applied across all modalities leads to synchronized evolution in time. To resolve this temporal inconsistency issue, we define a polynomial interpolation for position mainfold to explicitly decouples its evolution from other mainfolds. The interpolation is defined as:
$\mathbf{X}_{i}^{t}=\mathbf{X}_{i}^{0}+(1-(1-t)^{k})\cdot\mathbf{X}_{i}^{1},\quad t\in[0,1],$ (4)
@@
Non-Linear Flow and Diffusion Formulation.
The non-linear flow model is not introduced arbitrarily but is inspired by both flow matching and diffusion processes. In diffusion models, the distribution evolves over time via stochastic differential equations (SDEs), with the transformation given by $\mu_{t}(x)=\alpha_{1-t}x_{1}$ and the noise term $\sigma_{t}(x)=\sqrt{1-\alpha_{1-t}^{2}}$, where $\mu_{t}(x)$ represents the mean of the data distribution at time $t$ and $x_{1}$ is the target distribution.
It has also been shown that flow matching, as defined by the vector field, can be derived into the form of $\mu_{t}(x)$. The optimal transport (OT) assumption corresponds to a linear form: $\mu_{t}(x)=tx_{1}$, representing uniform evolution from the initial to the target distribution.
Building on this, our proposed non-linear flow corresponds to a time-varying interpolation: $\mu_{t}(x)=(1-(1-t)^{k})x_{1}$, which reduces to the linear flow when $k=1$. For $k>1$, the non-linear flow provides increased flexibility to simulate the varying convergence speed of the position manifold in docking tasks, accurately capturing the complex dynamics of peptide conformation changes.
Loss Function for Position Flow.
@@
Non-Uniform Time Sampling Strategy.
To address the bias of standard uniform sampling—where interpolated positions $\mathbf{X}_{t}$ cluster near $\mathbf{X}^{1}$ under polynomial-based interpolation—we redefine the time variable as $t=z^{k}$ ($z\sim\mathcal{U}(0,1)$). This non-uniform sampling adjusts the temporal density, prioritizing the initial phase ($t\to 0$) where rapid positional convergence occurs. By reshaping the probability distribution of $t$, the strategy ensures balanced coverage of the entire trajectory, preventing under-sampling of critical early dynamics. Simultaneously, it allocates more training samples to the information-sparse initial phase, enabling the model to robustly learn fast positional alignment governed by large initial gradients, while stabilizing gradient updates across all stages.
3.3 Linear Flow for Orientation and Amino Acid Type
In this section, we follow recent advances in flow-based methods (Lin et al., 2024; Li et al., 2024), modeling the flows of rotation, torsion, and amino acid types using the optimal transport assumption, which enables efficient transitions between states. As the peptide approaches the pocket, these manifolds adjust according to the pocket environment, ensuring temporal consistency. This adjustment allows us to treat these three modalities as time-synchronized linear flows. Together with the non-linear flow model discussed in the previous section, our approach decouples the evolution of the position manifold from the other manifolds, effectively capturing cross-modal temporal inconsistencies.
Rotation Matrices.

### v2

# THFlow: A Temporally Hierarchical Flow Matching Framework for 3D Peptide Design

@@ Abstract
Deep generative models provide a promising approach to de novo 3D peptide design. Most of them jointly model the distributions of peptide’s position, orientation, and conformation, attempting to simultaneously converge to the target pocket. However, in the early stage of docking, optimizing conformation-only modalities such as rotation and torsion can be physically meaningless, as the peptide is initialized far from the protein pocket and no interaction field is present. We define this problem as the multimodal temporal inconsistency problem and claim it is a key factor contributing to low binding affinity in generated peptides. To address this challenge, we propose THFlow, a novel flow matching-based multimodal generative model that explicitly models the temporal hierarchy between peptide position and conformation. It employs a polynomial-based conditional flow to accelerate positional convergence early on, and later aligns it with rotation and torsion for coordinated conformation refinement under the emerging interaction field. Additionally, we incorporate interaction-related features, such as polarity, to further enhance the model’s understanding of peptide-protein binding. Extensive experiments demonstrate that THFlow outperforms existing methods in generating peptides with superior stability, affinity, and diversity, offering an effective and accurate solution for advancing peptide-based therapeutic development.
I Introduction
@@
Peptides are short, single-chain proteins composed of amino acid. Therapeutic peptide plays a pivotal role in applications, and their design is essential for achieving precise and stable docking with target receptor proteins [1]. The rapid development of deep learning has made 3D peptide design a feasible computational task, offering a powerful approach for developing therapeutics against protein-related diseases [2, 3]. Recent generative models tackle the vast and complex chemical space by integrating protein sequence and structure data to extract rich feature representations, facilitating the exploration of sequence-structure relationships [4]. More recently, dynamic models such as diffusion and flow matching have gained prominence, simulating the physical docking process by generating trajectories for a peptide’s movement via time-dependent vector fields [5, 6]. These models explicitly operate on key physical modalities, including displacement, rotation, and other conformational degrees of freedom. By directly modeling the evolution of these dynamics, they enable efficient exploration of conformational space and provide a realistic simulation of the docking process. Thus, they overcome the limitations of traditional experimental methods, such as high costs and long timelines [7].
However, existing diffusion- and flow-based models struggle to generate peptides with both high binding affinity and structural stability [8]. One key limitation is their inability to model the temporal hierarchy across different modalities. Although these models generate multimodal variables over discrete timesteps, they often assume that all modalities evolve in a strictly synchronized fashion [9, 10, 11]. In reality, peptides must first move toward the target binding pocket, and only then undergo fine-grained conformational adjustments, such as backbone rotation and side-chain torsion, in response to the local interaction environment [12]. As a result, at each simulation timestep, the peptide’s position, conformation, and amino acid type are combined in a temporally misaligned manner, violating the natural progression of peptide dynamics. This multimodal temporal inconsistency problem leads to inaccurate modeling of peptide dynamics and reduces the biological plausibility of the simulation.
Figure 1: Temporally hierarchical peptide docking process
@@
To address this multimodal temporal inconsistency problem, we propose THFlow, a novel Temporally Hierarchical multimodal flow matching method. It accelerates the peptide’s movement toward the target binding pocket, followed by fine-tuning its conformation and posture, closely mirroring the actual docking process. To realize this, the core idea is to design a time-varying flow for the position modality, where the flow’s gradient vector decreases in magnitude over time, accelerating convergence in the initial stage. Meanwhile, the other modalities converge at a constant rate, effectively decoupling the evolution of position from conformation and posture, thus capturing the temporally hierarchical nature of the different modalities. Additionally, to prevent the model from focusing solely on conformational alignment while neglecting docking interactions, interaction-related information such as polarity, charge, and hydrophilicity is incorporated, ensuring stronger binding affinity to the target protein.
@@
The THFlow framework consists of two key components: (i) During the pre-training phase, the model learns the evolution of four modalities: peptide position, rotation, torsion, and amino acid types, from noise to the target distribution. A neural network is used to independently learn the gradient dynamics for each modality. (ii) In the sampling phase, based on ordinary differential equations (ODEs) [13], the pre-trained model predicts the gradient vector for the next timestep based on the current timestep’s information, progressively simulating the peptide docking process.
THFlow provides two key advantages: (i) it introduces a time-varying flow to decouple the convergence of different modalities, transforming the evolution from synchronized to asynchronous across temporal hierarchies. This enables more accurate peptide docking simulations, resulting in the generation of twice as many stable peptide-protein complexes compared to the baseline; (ii) the rapid convergence of the position modality reduces ineffective time spent on conformation exploration in the early stages, allowing the model to fine-tune conformation and posture near the binding pocket. By incorporating interaction-related information, the model explores interaction forces in greater depth, leading to a 7.13% increase in the affinity of the generated peptides.
@@
  • We first identify the multimodal temporal inconsistency problem as a critical issue in peptide design models, which violates biological principles and hinders the generation of high-affinity peptides.
@@
  • We design a novel temporally hierarchical flow matching framework THFlow that decouples the evolution of different modalities, providing a more accurate simulation of docking process.
@@
  • Extensive experiments demonstrate THFlow’s superior performance in both co-design and re-docking, significantly improving generated peptides’ stability and affinity, while providing an efficient method for advancing peptide-based therapeutic development.
II Related Work
II-A Peptide generation
@@
Peptide generation is a complex task in drug discovery, aiming to design peptides with specific biological functions, such as binding to target proteins or modulating enzymatic activity. Generative models have significantly advanced peptide design by exploring large sequence and structural spaces with greater flexibility. For example, AMP-diffusion [14] utilizes the capabilities of protein large language model ESM-2 [15] to regenerate functional antimicrobial peptides (AMPs). MMCD [16] employs multi-modal contrastive learning in diffusion, exploiting the integration of both sequence and structural information to produce peptides with high functional relevance. Besides, RFDiffusion [6], created for protein design, inspired the application of similar diffusion techniques for peptide generation.
@@
Peptide design is a subtask of peptide generation, involving several key components. One important aspect is backbone design, where methods like PepFlow [10] and PPFlow [11] use flow matching to simulate dynamic conformational changes and optimize peptide properties. Another aspect is side-chain packing, with methods such as RED-PPI [17] focusing on protein-protein complexes and DiffPack [18] targeting peptide-protein interactions.
II-B Protein–ligand docking
@@
Protein–ligand docking aims to predict the binding pose and affinity of a ligand to a target protein. Traditional docking methods often rely on rigid-body simulations and predefined scoring functions, which can struggle to handle flexible ligands and complex interactions [19]. Recent years, deep learning (DL) approaches have significantly improved docking accuracy. Models like AtomNet [20] and OnionNet [21] use convolutional neural networks (CNNs) to capture complex molecular features, significantly enhancing binding affinity predictions. Further developments have incorporated graph-based models, such as GraphSite [22] and DGraphDTA [23], which represent protein-ligand interactions as graphs to better model flexible docking. Additionally, advances in protein structure characterization, notably via AlphaFold [24], have enhanced understanding of protein flexibility, fueling the development of structure-based docking prediction methods [25]. However, challenges still remain in fully capturing the dynamic and flexible nature of ligands, highlighting the need for further improvements in docking models.
III Method
III-A Preliminary
III-A1 Problem Statement
@@
A peptide is a short sequence of amino acids, where each amino acid has a specific type, denoted by $a$, and consists of a rigid backbone and a flexible side chain.
@@
The backbone consists of four heavy atoms $N$, $_{\alpha}, CC$, and $O$, with the position of $C_{\alpha}$ serving as the reference, denoted as $\mathbf{X}_{i}$ at the origin. Attached to this alpha-carbon $C_{\alpha}$ is the side chain, a variable chemical group that determines its unique properties. Peptide’s orientation is modeled in two parts: the overall rigid rotation, represented by the rotation matrix $\mathbf{R}\in\operatorname{SO}(3)$, and the side-chain torsion, captures four torsion angles $\boldsymbol{\tau}=(\chi_{1},\chi_{2},\chi_{3},\chi_{4})$ of the rotatable bonds.
@@
III-A2 Conditional Flow Matching
Conditional Flow Matching (CFM) [26] models the transformation between a prior distribution $p=p_{0}$ and a target data distribution $q=p_{1}$ on a manifold $\mathcal{M}$ using ODEs. Mathematically, this “flow” represents a probability path generated by a time-dependent vector field. The transformation is described as:
$\frac{d}{dt}\varphi_{t}(x)=u_{t}(x),$ (2)
where $\varphi_{t}(x)$ represents the data distribution at time $t$, and $u_{t}(x)$ is the time-dependent vector field that driving the flow. In the flow matching framework, the true vector field $u_{t}(x)$ is considered intractable. Therefore, the method defines a conditional vector field $u_{t}(x|x_{1})$ based on a target data point $x_{1}$ to serve as a tractable learning target. A neural network is then trained to approximate this conditional field, and its output is denoted as the learned vector field $v_{t}(x)$. The model is trained to minimize the difference between its predicted vector field $v_{t}(x)$ and the conditional target field $u_{t}(x|x_{1})$ via the CFM loss:
In most current methods, the distribution $\varphi_{t}(x)$ is typically assumed to follow a simple linear interpolation between the initial distribution $x_{0}$ and target distribution $x_{1}$: $\varphi_{t}(x)=tx_{1}+(1-t)x_{0}$, where $t\in[0,1]$. The corresponding vector field is $u_{t}(x|x_{1})=\frac{d}{dt}\varphi_{t}(x)=x_{1}-x_{0}$. This form has been proven to correspond to the Optimal Transport (OT) solution [26], which provides an efficient and fast means of transforming distributions with a fixed direction and magnitude.
III-A3 Time-varying Flow and Diffusion Formulation
Both diffusion and flow matching model the transformation of distributions as an evolving path $\varphi_{t}(x)$, which can be defined as an interpolation of the target signal and the initial noise:
$\varphi_{t}(x)=\mu(t)x_{1}+\sigma(t)x_{0}.$ (4)
Here, $\mu(t)$ and $\sigma(t)$ are time-dependent scalar coefficients satisfying the boundary conditions $\mu(0)=0,\sigma(0)=1$ and $\mu(1)=1,\sigma(1)=0$.
In diffusion models, the transformation is typically defined by variance-preserving stochastic differential equations (SDEs) [27]. This corresponds to a constraint where the sum of squares of the coefficients is one: $\mu(t)^{2}+\sigma(t)^{2}=1$. This formulation maintains signal stability by ensuring that the variance of the evolving distribution remains constant throughout the transformation.
In contrast, flow matching under the OT assumption defines a direct, linear path. It is realized with $\mu(t)=t$ and $\sigma(t)=1-t$, which corresponds to a simpler linear constraint where the coefficients sum to one: $\mu(t)+\sigma(t)=1$. This linear structure provides an intuitive and efficient straight-line trajectory from noise to data.
Building on this, our proposed time-varying flow corresponds to a polynomial interpolation: $\mu(t)=1-(1-t)^{k}$ and $\sigma(t)=(1-t)^{k}$, which also adheres to the linear sum constraint $\mu(t)+\sigma(t)=1$, and reduces to the linear flow when $k=1$. For $k>1$, the flow provides increased flexibility to simulate the varying convergence speed of the position manifold in docking tasks, accurately capturing the complex dynamics of peptide conformation changes.
III-B Time-varying Flow for Position
@@
Despite being efficient and fast, the OT assumption applied across all modalities leads to synchronized evolution in time. To resolve this multimodal temporal inconsistency problem, we define a polynomial interpolation for position mainfold to explicitly decouples its evolution from other mainfolds. The interpolation is defined as:
$\mathbf{X}_{i}^{t}=(1-t)^{k}\mathbf{X}_{i}^{0}+(1-(1-t)^{k})\cdot\mathbf{X}_{i}^{1},\quad t\in[0,1],$ (5)
@@
Loss Function for Position Flow
@@
Figure 3: Overview of the THFlow framework
Non-Uniform Time Sampling Strategy
To address the bias of standard uniform sampling—where interpolated positions $\mathbf{X}^{t}$ cluster near $\mathbf{X}^{1}$ under polynomial-based interpolation—we redefine the time variable as: $t=z^{k},z\sim\mathcal{U}(0,1)$. This non-uniform sampling adjusts the temporal density, prioritizing the initial phase ($t\to 0$) where rapid positional convergence occurs. By reshaping the probability distribution of $t$, the strategy ensures balanced coverage of the entire trajectory, preventing under-sampling of critical early dynamics. Simultaneously, it allocates more training samples to the information-sparse initial phase, enabling the model to robustly learn fast positional alignment governed by large initial gradients, while stabilizing gradient updates across all stages.
III-C Linear Flow for Orientation and Amino Acid Type
In this section, we follow recent advances in flow-based methods [11, 10], modeling the flows of rotation, torsion, and amino acid types using the optimal transport assumption, which enables efficient transitions between states. As the peptide approaches the pocket, these manifolds adjust according to the pocket environment, ensuring temporal consistency. This adjustment allows us to treat these three modalities as time-synchronized linear flows. Together with the time-varying flow model discussed in the previous section, our approach decouples the evolution of the position manifold from the other manifolds, effectively capturing cross-modal temporal inconsistencies.
III-C1 Rotation Matrices